ALGEBRAIC INVARIANTS FOR RIGHT-ANGLED ARTIN GROUPS

STEFAN PAPADIMA ALEXANDER I. SUCIU

Abstract. A finite simplicial graph Γ determines a right-angled Artin group G_{Γ}, with generators corresponding to the vertices of Γ, and with a relation vw = wv for each pair of adjacent vertices. We compute the lower central series quotients, the Chen quotients, and the (first) resonance variety of G_{Γ}, directly from the graph Γ.